OpenStudy (sh3lsh):

A mass m sliding on a frictionless surface and connected to a spring with a positive, nonzero spring constant follows the DE mx'' + kx = 0. Note F = -kx. The energy of a solution to this equation is defined by E(t): = 1/2 (x')^2 + (1/2)kx^2. Show E(t) is constant.

OpenStudy (sh3lsh):

Perhaps a nudge in the right direction is the best. I'm unsure of how to begin.

OpenStudy (roadjester):

This is a physics question isn't it?

OpenStudy (sh3lsh):

Eh, it's a Differential Equation question.

OpenStudy (roadjester):

Is this your givens: \[{\frac {d^2m} {dx^2} +kx=0}\] \[E(t)={\frac 1 2}({\frac {d^2m}{dx^2})^2 + {\frac 1 2kx^2}}\]

OpenStudy (sh3lsh):

I made a mistake. I'm sorry. E(t) is supposed to be the first derivative of x, not second. \[m \frac{ d ^{2x} }{ dt ^{2} } + kx = 0\] \[\frac{ 1 }{ 2 } m (\frac{ dx }{ dt })^{2} + \frac{ 1 }{ 2 } kx^{2}\]

OpenStudy (sh3lsh):

Sorry again, \[m \frac{ d ^{2}x }{ d t^{2} }\]

OpenStudy (sh3lsh):

It also notes that the first equation is really F=ma since F exerted on the mass by the spring is -kx

OpenStudy (roadjester):

Is the E(t) equation correct? I seems a bit odd to square a second derivative

OpenStudy (sh3lsh):

Equations: \[m \frac{ d^{2}x }{ dt ^{2} } + kx = 0\] \[E(t) := \frac{ 1 }{ 2 } m (\frac{ dx }{ dt })^{2} + \frac{ 1 }{ 2 } kx^{2}\]

OpenStudy (sh3lsh):

Would you have any idea on how to approach this problem?

OpenStudy (roadjester):

It's been a while since i've done DE, so maybe not, but give me a sec

OpenStudy (roadjester):

Not sure if this will help, but take a look at page 1. http://www.stewartcalculus.com/data/CALCULUS%20Concepts%20and%20Contexts/upfiles/3c3-AppsOf2ndOrders_Stu.pdf What I'm thinking is, if you obtain the arbitrary solution, take the first derivative, and substitute it into your E(t) equation, everything should cancel (i think) if E(t) is a constant. (Not too confident on that last part)

OpenStudy (irishboy123):

you have: \(m \ddot x + k x = 0 \) which is same as: \(m \dot x \frac{d\dot x}{dx} + k x = 0\)...continuity assumed. and from \(m \dot x \frac{d\dot x}{dx} + k x = 0\), we have \(\frac{1}{2} m \dot x ^ 2 + \frac{1}{2} k x ^ 2 = const \) and you are choosing to equate that expression with E(t)....... if that is true, then \(\dot E = \dot {const} = 0\)

OpenStudy (irishboy123):

this is all very circular :-) but very interesting also.

OpenStudy (sh3lsh):

Thanks you all!